If the price of suntan lotion increases from $6 to $8 per bottle and quantity demanded decrease from 900,000 bottles to 845,000 bottles, what is the price elasticity of demand for suntan lotion?
Answer #1

Price elasticity of demand = % change in quantity demanded/%change in price

% change in quantity demanded = ((845000-900000)/900000)*100 = -6.11%

% change in price = ((8-6)/6)*100 = 33.33%

So, PED = -6.11/33.33 = 0.18

Hence the price elasticity of demand for suntan lotion is 0.18, means the demand is inelastic for price changes.